# Tri-resonant leptogenesis from modular symmetry neutrino models

### v1

@@ Abstract
In this paper, we have studied the consequences of some representative modular symmetry neutrino models for tri-resonant leptogenesis (which is realized by having three nearly degenerate right-handed neutrinos). To be specific, we have considered modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix $M_{\rm R}$ as shown in Eq. (8) which gives three degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work. For these models, we have considered two minimal ways to generate the desired small mass splittings among the three right-handed neutrinos: one scenario is to modify $M_{\rm R}$ to a form as shown in Eq. (10), while another scenario is to consider the renormalization-group corrections for the right-handed neutrino masses. In the former scenario, for the considered models, the observed value of $\eta_{\rm B}$ can always be successfully reproduced for appropriate values of $\mu$. But in the latter scenario not all the considered models can successfully reproduce the observed value of $\eta_{\rm B}$.
@@ 1 Introduction
As we know, the discovery of neutrino oscillations reveals that neutrinos are massive and their flavor eigenstates $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) are certain superpositions of their mass eigenstates $\nu_{i}$ (for $i=1,2,3$) possessing definite masses $m_{i}$: $\nu_{\alpha}=\sum_{i}U_{\alpha i}\nu_{i}$ with $U_{\alpha i}$ being the $\alpha i$-th element of the neutrino mixing matrix $U$ [1]. In the standard form, $U$ is expressed in terms of three mixing angles $\theta_{ij}$ (for $ij=12,13,23$), one Dirac CP phase $\delta$ and two Majorana CP phases $\alpha_{21}$ and $\alpha_{31}$ as
@@
Nowadays, the neutrino mixing angles and neutrino mass squared differences $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$ have been measured to good degrees of accuracy, and there is also a preliminary result for $\delta$ (but with a large uncertainty). Several research groups have performed global analyses of the existing neutrino oscillation data to extract the values of these parameters [2, 3]. For reference, the results in Ref. [2] are reproduced in Table 1 here. Note that the sign of $\Delta m^{2}_{31}$ remains undetermined, thereby allowing for two possible neutrino mass orderings: the normal ordering (NO) $m_{1}<m_{2}<m_{3}$ and inverted ordering (IO) $m_{3}<m_{1}<m_{2}$. However, neutrino oscillations are completely insensitive to the absolute neutrino mass scale and the Majorana CP phases. Their values can only be inferred from certain non-oscillatory experiments such as the neutrinoless double beta decay experiments [4]. But so far there has not been any lower bound on the lightest neutrino mass, nor any constraint on the Majorana CP phases.
For the neutrino mixing angles, Table 1 shows that $\theta_{12}$ and $\theta_{23}$ are close to some special values (i.e., $\sin^{2}\theta_{12}\sim 1/3$ and $\sin^{2}\theta_{23}\sim 1/2$). But $\theta_{13}$ is relatively small. In fact, before its value was measured, $\theta_{13}$ had been widely expected to be vanishingly small. For the ideal case of $\sin^{2}\theta_{12}=1/3$, $\sin^{2}\theta_{23}=1/2$ and $\theta_{13}=0$, the neutrino mixing matrix takes a very simple form as
@@
which is referred to as the tribimaximal (TBM) mixing [5]. This particular mixing has inspired intensive model-building studies with the employment of certain discrete non-Abelian symmetries (e.g., ${\rm A}_{4}$ and ${\rm S}_{4}$) [6].
@@
Noteworthy, a shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping. Moreover, higher-dimensional operators with more flavon insertions and unknown coefficients are often present, affecting the predictive power of the models. In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular [7, 8]. In the modular-invariance approach, the flavour symmetry is broken by the VEV of a single scalar field—the so-called modulus, while the flavon fields are not necessary. And the Yukawa couplings and fermion mass terms are certain special functions of the modulus called modular forms that have particular properties under the modular group transformations. Moreover, all higher-dimensional operators are unambiguously determined in the limit of unbroken supersymmetry. As a result, flavor models with modular invariance depend on fewer parameters, enhancing their predictive power.
@@
As for the neutrino masses, one of the most popular and natural ways of generating them is the type-I seesaw model [9] in which three heavy right-handed neutrinos $N_{I}$ ($I=1,2,3$) are introduced into the Standard Model (SM). First of all, $N_{I}$ can constitute the Yukawa coupling operators together with the lepton doublets $L_{\alpha}$ and the Higgs doublet $H$: $Y_{I\alpha}\overline{N_{I}}HL_{\alpha}$ with $Y_{I\alpha}$ being the $I\alpha$-th element of the Yukawa coupling matrix $Y$. These operators will generate the Dirac neutrino mass terms $(M_{\rm D})_{I\alpha}=Y_{I\alpha}v$ [here $(M_{\rm D})_{I\alpha}$ is the $I\alpha$-th element of the Dirac neutrino mass matrix $M_{\rm D}$] after the neutral component of $H$ acquires a nonzero vacuum expectation value (VEV) $v=174$ GeV. In addition, $N_{I}$ themselves can also have the Majorana mass terms $\overline{N_{I}}(M_{\rm R})_{IJ}N^{c}_{J}$ [here $(M_{\rm R})_{IJ}$ are the $IJ$-th element of the right-handed neutrino mass matrix $M_{\rm R}$]. Then, under the seesaw condition $M_{\rm R}\gg M_{\rm D}$, integrating the right-handed neutrinos out yields an effective Majorana mass matrix for three light neutrinos as
@@
Here we would like to point out that throughout the paper the fermion mass matrices are written in the right-left convention which is more conveniently used in the supersymmetric setup as applicable to our study.
Normal Ordering Inverted Ordering
bfp ±$\pm 1\sigma$ $3\sigma$ range bfp ±$\pm 1\sigma$ $3\sigma$ range
$\sin^{2}\theta_{12}$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$ $0.303_{-0.011}^{+0.012}$ $0.270\to 0.341$
$\sin^{2}\theta_{23}$ $0.572_{-0.023}^{+0.018}$ $0.406\to 0.620$ $0.578_{-0.021}^{+0.016}$ $0.412\to 0.623$
$\sin^{2}\theta_{13}$ $0.02203_{-0.00059}^{+0.00056}$ $0.02029\to 0.02391$ $0.02219_{-0.00057}^{+0.00060}$ $0.02047\to 0.02396$
$\delta/\pi$ $1.09_{-0.14}^{+0.23}$ $0.96\to 1.33$ $1.59_{-0.18}^{+0.15}$ $1.41\to 1.74$
$\Delta m^{2}_{21}/(10^{-5}~{}{\rm eV}^{2})$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$ $7.41_{-0.20}^{+0.21}$ $6.82\to 8.03$
$\Delta m^{2}_{3\ell}/(10^{-3}~{}{\rm eV}^{2})$ $2.511_{-0.027}^{+0.028}$ $2.428\to 2.597$ $-2.498_{-0.025}^{+0.032}$ $-2.581\to-2.408$

### v3

# Renormalization group running triggered tri-resonant leptogenesis within neutrino flavor-symmetry models

@@ Abstract
In this paper, we have studied the consequences of some representative neutrino flavor-symmetry models for tri-resonant leptogenesis (which is realized by having three nearly degenerate right-handed neutrinos). To be specific, we have considered a neutrino flavor-symmetry model realizing the TM1 mixing and modular ${\rm A}_{4}$, ${\rm S}_{4}$ and ${\rm A}_{5}$ symmetry models that have a right-handed neutrino mass matrix $M_{\rm R}$ as shown in Eq. (15) which gives three exactly degenerate right-handed neutrino masses and consequently prohibits the leptogenesis mechanism to work successfully. For these models, we study the scenario that the desired right-handed neutrino mass splittings for leptogenesis to work are generated from the renormalization-group running effects. In such a scenario, we explore the parameter space that allows for the reproduction of the observed baryon-antibaryon asymmetry of the Universe.
@@ 1 Introduction
As we know, the discovery of neutrino oscillations reveals that neutrinos are massive and their flavor eigenstates $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) are certain superpositions of their mass eigenstates $\nu_{i}$ (for $i=1,2,3$) possessing definite masses $m_{i}$: $\nu_{\alpha}=\sum_{i}U_{\alpha i}\nu_{i}$ with $U_{\alpha i}$ being the $\alpha i$-element of the neutrino mixing matrix $U$ [1]. In the standard form, $U$ is expressed in terms of three mixing angles $\theta_{ij}$ (for $ij=12,13,23$), one Dirac CP phase $\delta$ and two Majorana CP phases $\alpha_{21}$ and $\alpha_{31}$ as
@@
Nowadays, the neutrino mixing angles and neutrino mass squared differences $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$ have been measured to good degrees of accuracy, and there is also a preliminary result for $\delta$ (but with a large uncertainty). Several research groups have performed global analyses of the existing neutrino oscillation data to extract the values of these parameters [2]. As is known, the sign of $\Delta m^{2}_{31}$ remains undetermined, thereby allowing for two possible neutrino mass orderings: the normal ordering (NO) $m_{1}<m_{2}<m_{3}$ and inverted ordering (IO) $m_{3}<m_{1}<m_{2}$. On the other hand, neutrino oscillations are completely insensitive to the absolute neutrino mass scale and the Majorana CP phases. Their values can only be inferred from certain non-oscillatory experiments such as the neutrinoless double beta decay experiments [3]. But so far there has not been any lower bound on the lightest neutrino mass, nor any constraint on the Majorana CP phases.
For the neutrino mixing angles, it is known that $\theta_{12}$ and $\theta_{23}$ are close to some special values (i.e., $\sin^{2}\theta_{12}\sim 1/3$ and $\sin^{2}\theta_{23}\sim 1/2$), while $\theta_{13}$ is relatively small. In fact, before the true value of $\theta_{13}$ was experimentally pinned down, it had been widely believed to be vanishingly small. For the ideal case of $\sin^{2}\theta_{12}=1/3$, $\sin^{2}\theta_{23}=1/2$ and $\theta_{13}=0$, the neutrino mixing matrix takes a very simple form as
@@
which is referred to as the tribimaximal (TBM) mixing [4]. This particular mixing has inspired intensive model-building studies with the employment of certain discrete non-Abelian symmetries (e.g., ${\rm A}_{4}$ and ${\rm S}_{4}$) [5]. However, the observation of a relatively large $\theta_{13}$ (compared to 0) motivates us to make corrections for the TBM mixing. In this connection, an economical and natural choice is to retain its first or second column while modifying the other two columns within the unitary constraints, thus yielding the first or second trimaximal (TM1 or TM2) mixing [6]
$U_{\rm TM1}=\displaystyle\frac{1}{\sqrt{6}}\left(\begin{array}[]{ccc}2&\cdot&\cdot\cr 1&\cdot&\cdot\cr 1&\cdot&\cdot\cr\end{array}\right)\;,\hskip 28.45274ptU_{\rm TM2}=\displaystyle\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}\cdot&1&\cdot\cr\cdot&-1&\cdot\cr\cdot&-1&\cdot\cr\end{array}\right)\;,$ (12)
where the dot signs denote the unspecified elements. Due to their simplicity and interesting consequences, the TM1 and TM2 mixings have attracted a lot of attention and many neutrino flavor-symmetry models have been formulated to realize them [5].
@@
An obvious shortcoming of the conventional flavor-symmetry models is that many flavon fields need to be introduced in order to achieve a desired breaking pattern of the flavor symmetry, and the scalar potential of these flavon fields needs a complicated symmetric shaping. Moreover, higher-dimensional operators with more flavon insertions and unknown coefficients are often present, affecting the predictive power of the models. In recent years, the idea of modular invariance which provides substantial simplifications to these problems has been proposed and become increasingly popular [7, 8]. In the modular-invariance approach, the flavor symmetry is broken by the vacuum expectation value (VEV) of a single scalar field—the so-called modulus, while the flavon fields are not necessary. And the Yukawa couplings and fermion mass terms are certain special functions of the modulus called modular forms that have particular properties under the modular group transformations. Moreover, all higher-dimensional operators are unambiguously determined in the limit of unbroken supersymmetry. As a result, flavor models with modular invariance depend on fewer parameters, enhancing their predictive power.
@@
As for the neutrino masses, one of the most popular and natural ways of generating them is the type-I seesaw model [9] in which three heavy right-handed neutrinos $N_{I}$ ($I=1,2,3$) are introduced into the Standard Model (SM). First of all, $N_{I}$ can constitute the Yukawa coupling operators together with the lepton doublets $L_{\alpha}$ and the Higgs doublet $H$: $Y_{I\alpha}\overline{N_{I}}HL_{\alpha}$ with $Y_{I\alpha}$ being the $I\alpha$-element of the Yukawa coupling matrix $Y$. These operators will generate the Dirac neutrino mass terms $(M_{\rm D})_{I\alpha}=Y_{I\alpha}v$ (here $(M_{\rm D})_{I\alpha}$ is the $I\alpha$-element of the Dirac neutrino mass matrix $M_{\rm D}$) after the neutral component of $H$ acquires a nonzero VEV $v=174$ GeV. In addition, $N_{I}$ themselves can also have the Majorana mass terms $\overline{N_{I}}(M_{\rm R})_{IJ}N^{c}_{J}$ (here $(M_{\rm R})_{IJ}$ is the $IJ$-element of the right-handed neutrino mass matrix $M_{\rm R}$). Then, under the seesaw condition $M_{\rm R}\gg M_{\rm D}$, integrating the right-handed neutrinos out yields an effective Majorana mass matrix for three light neutrinos as
@@
Here we would like to emphasize that throughout the paper the fermion mass matrices are written in the right-left convention which is more conveniently used in the supersymmetric setup as applicable to our study.
Remarkably, the seesaw model also provides an attractive explanation (known as the leptogenesis mechanism [10, 11]) for the baryon-antibaryon asymmetry of the Universe [12]
$\eta_{\rm B}\equiv\frac{n_{\rm B}-n_{\rm\bar{B}}}{n_{\gamma}}\simeq(6.11\pm 0.04)\times 10^{-10}\;,$ (14)
where $n_{\rm B}$ ($n_{\rm\bar{B}}$) denotes the baryon (antibaryon) number density and $n_{\gamma}$ the photon number density. The leptogenesis mechanism works in a way as follows: a lepton-antilepton asymmetry $\eta_{\rm L}\equiv(n_{\rm L}-n_{\rm\bar{L}})/n_{\gamma}$ is firstly generated from the out-of-equilibrium and CP-violating decays of the right-handed neutrinos and then partly converted into the baryon-antibaryon asymmetry via the sphaleron processes: $\eta_{\rm B}\simeq-c\eta_{\rm L}$ with $c=28/79$ or $8/23$ in the SM or MSSM (Minimal Supersymmetric Standard Model) framework.
Unfortunately, for some typical neutrino flavor-symmetry models (see those given in sections 3-6), the right-handed neutrino mass matrix takes a form as
$M_{\rm R}=M_{0}\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\;$, (15)
which gives three exactly degenerate right-handed neutrino masses (i.e., $M_{0}$). In this case, the leptogenesis mechanism is prohibited to work successfully (to be specific, there would no CP asymmetries for the decays of right-handed neutrinos). But just as a coin has two sides, this kind of models can serve as a unique basis for realizations of the tri-resonant leptogenesis scenario [13]: if the right-handed neutrino masses receive certain corrections so that their degeneracy is lifted to an appropriate extent, then the tri-resonant leptogenesis scenario will be naturally realized.
In this paper, within this kind of neutrino flavor-symmetry models, we study the scenario that the desired right-handed neutrino mass splittings for leptogenesis to work are generated from the renormalization-group equation (RGE) effects [14]: if the energy scale $\Lambda_{\rm FS}$ of the flavor-symmetry physics that shapes the special textures of neutrino mass matrices is much higher than the right-handed neutrino mass scale $M_{0}$ (throughout this paper, $M_{0}$ will be used to denote the common masses of three nearly degenerate right-handed neutrinos when their splitting is not important) where leptogenesis takes place, then the RGE effect between the scales of $\Lambda_{\rm FS}$ and $M_{0}$ may induce some mass splittings among the three right-handed neutrinos (see section 2.2). Such a correction of $M_{\rm R}$ has the following two merits: 1) this effect is spontaneous, provided that there is a considerable gap between the flavor-symmetry scale and the right-handed neutrino mass scale; 2) this effect is minimal, in the sense that it does not need to introduce additional parameters.
The remaining parts of this paper are organized as follows. In section 2 we will first give some basic formulas about leptogenesis and RGE effects that will be used in our study. Then, we perform the study for four representative neutrino flavor-symmetry models which contain three degenerate right-handed neutrinos For a similar study about the $\Delta(3n^{2})$ and $\Delta(6n^{2})$ neutrino flavor-symmetry models, see Ref. [15].. We successively study a neutrino flavor-symmetry model realizing the TM1 mixing in section 3, modular ${\rm A}_{4}$ symmetry model in section 4, modular ${\rm S}_{4}$ symmetry model in section 5, and modular ${\rm A}_{5}$ symmetry model in section 6. In section 7, we summarize our main results.
2 Some formulas about leptogenesis and RGE effects
2.1 Some formulas about leptogenesis
In the resonant leptogenesis scenario, the CP asymmetries for the decay processes of right-handed neutrinos will be greatly enhanced due to nearly degenerate masses of the right-handed neutrinos. The resonance effects can be encoded in the effective neutrino Yukawa couplings as [13]
$\hskip-56.9055pt\left(\bar{Y}_{+}\right)_{I\alpha}=Y_{I\alpha}-{\rm i}V_{I\alpha}+{\rm i}\sum_{J,K=1}^{3}\left|\epsilon_{IJK}\right|Y_{J\alpha}$
$\hskip-5.69046pt\times\frac{M_{I}\left(\mathcal{M}_{IIJ}+\mathcal{M}_{JJI}\right)+{\rm i}R_{IK}\left[\mathcal{M}_{IKJ}\left(\mathcal{M}_{IIK}+\mathcal{M}_{KKI}\right)+\mathcal{M}_{JJK}\left(\mathcal{M}_{IKI}+\mathcal{M}_{KIK}\right)\right]}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}-2{\rm i}{\rm Im}(R_{IK})\left[M^{2}_{I}\left|A_{JK}\right|^{2}+M_{J}M_{K}{\rm Re}(A^{2}_{JK})\right]}\;,$ (16)
$A_{IJ}=\frac{\left(YY^{\dagger}\right)^{*}_{IJ}}{16\pi}\;,\hskip 28.45274ptV_{I\alpha}=\sum_{\beta=e,\mu,\tau}\sum_{K\neq I}\frac{Y^{*}_{I\beta}Y_{K\beta}Y_{K\alpha}}{16\pi}f\left(\frac{M^{2}_{K}}{M^{2}_{I}}\right)\;,$
$R_{IJ}=\frac{M^{2}_{I}}{M^{2}_{I}-M^{2}_{J}-2{\rm i}M^{2}_{I}A_{JJ}}\;,$ (17)
with $f(x)=\sqrt{x}[1-(1+x)\ln(1+1/x)]$ being the loop function. And the corresponding CP-conjugate effective Yukawa couplings $\left(\bar{Y}_{-}\right)_{I\alpha}$ can be obtained by simply making the replacement $Y_{I\alpha}\to Y^{*}_{I\alpha}$ in the above expressions. In terms of these effective Yukawa couplings, the decay widths of right-handed neutrinos can be expressed as
$\Gamma_{I}\equiv\sum_{\alpha}\Gamma\left(N_{I}\to L_{\alpha}+H\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}\right|^{2}\;,\quad\bar{\Gamma}_{I}\equiv\sum_{\alpha}\Gamma\left(N_{I}\to\overline{L}_{\alpha}+\overline{H}\right)=\frac{M_{I}}{16\pi}\left|\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}\right|^{2}\;,$ (18)
Finally, the CP-asymmetry matrix element $\varepsilon^{\left(I\right)}_{\alpha\beta}$ describing the asymmetries generated from the decays of $N_{I}$ can be obtained as
$\varepsilon^{\left(I\right)}_{\alpha\beta}=\frac{\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}\bar{\Gamma}_{I}-\mathcal{P}^{\left(I\right)}_{\alpha\beta}\Gamma_{I}}{\Gamma_{I}+\bar{\Gamma}_{I}}\;.$ (19)
Here the two projectors $\mathcal{P}^{\left(I\right)}_{\alpha\beta}$ and $\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}$ are given by
$\mathcal{P}^{\left(I\right)}_{\alpha\beta}\equiv C_{\alpha I}C^{*}_{\beta I}\;,\qquad\bar{\mathcal{P}}^{\left(I\right)}_{\alpha\beta}\equiv\bar{C}_{\alpha I}\bar{C}^{*}_{\beta I}\;,$ (20)
with
$C_{I\alpha}=\frac{\left(\bar{Y}_{+}\right)_{I\alpha}}{\sqrt{\left(\bar{Y}_{+}\bar{Y}^{\dagger}_{+}\right)_{II}}}\;,\qquad\bar{C}_{I\alpha}=\frac{\left(\bar{Y}_{-}\right)_{I\alpha}}{\sqrt{\left(\bar{Y}_{-}\bar{Y}^{\dagger}_{-}\right)_{II}}}\;.$ (21)
Neglecting CP-violating corrections, $C_{\alpha I}$ and $\bar{C}_{\alpha I}$ reduce to
$C^{0}_{I\alpha}=\frac{Y_{I\alpha}}{\sqrt{\left(YY^{\dagger}\right)_{II}}}\;.$ (22)